OpenStudy (anonymous):

1.) A park in a subdivision has a triangular shape. Two adjacent sides of the park are 533 and 525 feet. The angle between the sides is 53 degrees. Find the area of the park to the nearest square foot. My answer: 111,739. Is that correct? 2.) The diameter of a basketball rim is 18 inches. A standard basketball has a circumference of 30 inches. What is the distance between the ball and the rim in a shot in which the ball goes exactly in the center of the rim? My answer: 4.225. Is that correct?

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@ganeshie8 @Michele_Laino

OpenStudy (anonymous):

Oh wow this is worth a lot of point. I want to put every ounce of detail there can be.

OpenStudy (anonymous):

*points

OpenStudy (welshfella):

area = 0.5 ab sin C where a and b are the 2 sides and C is the included angle

OpenStudy (welshfella):

yes correct

OpenStudy (anonymous):

For both?

OpenStudy (anonymous):

There needs to be more detail I feel. What else should I put?

OpenStudy (welshfella):

no thats ok

OpenStudy (anonymous):

Its worth a ton of points, I think they need more information, that is what I am afraid of.

OpenStudy (welshfella):

just write the formula then the calculation and the answer

OpenStudy (welshfella):

2 work out the diameter of the basket ball first using circumference = pi* d then the requtred distance is (30 - d) / 2

OpenStudy (michele_laino):

for first question, I got the same result

OpenStudy (welshfella):

ooops sorry - i made a mistake in number 2

OpenStudy (michele_laino):

for the second question I got the same result

OpenStudy (welshfella):

the required value = ( 18 - d)/2 where d = diameter of the basketball and this comes to same as you got

OpenStudy (anonymous):

Alright thank you.

OpenStudy (anonymous):

So I still did it correctly

OpenStudy (rhr12):

You are correct for the second one

OpenStudy (rhr12):

and i gave u the answer for first one

OpenStudy (anonymous):

@ParthKohli can you give me more information for number 2?

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

anyone???

OpenStudy (triciaal):

not sure what you need your answers were verified

OpenStudy (anonymous):

The way it was solved.

OpenStudy (anonymous):

@welshfella can you tell me how number 2 was solved

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ParthKohli